2. A second enzyme, baaRysake, catalyzes the reaction of sad to happy. However, an inhibitor, anger, has been found that slows the rate of the reaction. the data below to answer the questions regarding this enzyme a. Make a rk plot for this data Make sure that there are osa, labels with units, legend and appropriate scale on the graph. Please remove the guidelines. Make sure the graph is large enough to read. IS pts) b. What type of enzymatic catalysis is taking place (based on you graph)? pts c. Determine the value for K and V 2 pts Angel mM 5.20 1996 16.83 13.11 13.33 21.41 18.15 14.29
Answer #1

Lineweaver -Burk equation-

1/V=(Km/Vmax)*1/[S] +1/Vmax

where V=reaction velocity

km=michaelis-menten constant

Vmax=maximum reaction velocity

[S]=substrate concentration

for 1/V vs 1/[S] plot ,a linear graph is obtained with slope=Km/Vmax, intercept=1/Vmax

[as per the linear eqn y=mx+c with m=slope,c=intercept]

b) it is competitive inhibition ,as km has increases ,but Vmax is unaffected ,intersects at 1/Vmax point(y-intercept same)

c) from the plot 1 slope=km/Vmax=0.071

from plot 2,slope=km'/Vmax=0.103

1/Vmax=y-intercept=0.046

Vmax=1/0.046=21.739 mM/s

plot 1,km=0.071*Vmax=1.543 mM/s

plot 2,km'=0.103*Vmax=2.239 mM/s